# Power Minimization for NOMA-assisted Pinching Antenna Systems With Multiple Waveguides

### v1

@@ II System Model
$\bm{x}_{n}=\sum_{m=1}^{M}\sqrt{P_{n}\alpha_{n,m}}\bm{x}_{n,m},$ (1)
where $\bm{x}_{n,m}$ and $\alpha_{n,m}$ denote the signal and the power allocation coefficient for user $m$ served by waveguide $n$, respectively.
@@
$h_{n,m}=\sum_{i=1}^{M}\frac{\sqrt{\eta}e^{-j\frac{2\pi}{\lambda}|\tilde{\psi}_{n,m}-\psi_{n,i}|}}{|\tilde{\psi}_{n,m}-\psi_{n,i}|}e^{-j\theta_{n,i}},$ (2)
in which $\eta=\frac{c^{2}}{16\pi^{2}f^{2}_{c}}$, $c$ is the speed of light, $f_{c}$ is the carrier frequency and $\theta_{n,i}=2\pi\frac{\psi_{n,0}-\psi_{n,i}}{\lambda_{g}}$ depicts the phase shift experienced at the $i$-th antenna on waveguide $n$. Thereof, $\psi_{n,0}$ and $\lambda_{g}$ are the location of the feed point of waveguide $n$ and the waveguide wavelength in a dielectric waveguide [5]. Moreover, we define $h_{n^{\prime},n,i,m}$ as the channel gain between the $i$-th antenna on the waveguide $n^{\prime}$ and the $m$-th user associated with the waveguide $n$. Mathematically, we have
$h_{n^{\prime},n,i,m}=\frac{\sqrt{\eta}e^{-j2\pi(\frac{1}{\lambda}|\tilde{\psi}_{n,m}-\psi_{n^{\prime},i}|+\frac{1}{\lambda_{g}}|\psi_{n^{\prime},0}-\psi_{n^{\prime},i}|)}}{|\tilde{\psi}_{n,m}-\psi_{n^{\prime},i}|}.$ (3)
@@
in which $\sigma^{2}$ represents the additive white Gaussian noise power. For the users served by the waveguide $n$, we rank their nSINRs in ascending order. Denote by $\bm{o}_{n}=(o_{1},o_{2},\ldots,o_{M})$, which satisfies $\text{nSIINR}_{n,o_{i}}<\text{nSIINR}_{n,o_{j}}$ if $i<j$. Following the downlink NOMA decoding principle, the interference plus noise ratio of user $o_{m}$ served by the waveguide $n$, denoted by $\text{SINR}_{n,o_{m}}$, can be expressed as follows:
@@ III Problem Formulation
In this paper, our objective is to minimize the total transmit power of all users while satisfying the data rate requirement of each user. For simplicity, we define $\bm{\alpha}_{n}=(\alpha_{n,1},\alpha_{n,2},\ldots,\alpha_{n,M})$ as the power allocation coefficients for users associated with the waveguide $n$. In addition, let $\bm{\alpha}=(\bm{\alpha}_{1},\bm{\alpha}_{2},\ldots,\bm{\alpha}_{N})$ be the power control coefficients for all users. We also denote $\bm{P}=(P_{1},P_{2},\ldots,P_{N})$ as the power allocation vector of all the waveguides. Furthermore, denote by $\bar{R}_{n,m}$ the minimum data rate requirement for user $m$ served by the waveguide $n$. Mathematically, we have
$\underset{\bm{\alpha},\bm{P}}{\text{min}}\sum_{n=1}^{N}P_{n}$ (8)
@@
where C1 indicates that, for each waveguide, the sum of the power coefficients of its served users should be 1. C2 ensures that the power allocation coefficient for each user lies between 0 and 1. C3 limits the maximum transmission power of waveguide $n$ to $\bar{P}_{n}$. Finally, C4 ensures that each user achieves its minimum data rate. The formulated optimization problem (8) is non-convex because $R_{n,m}$ is non-convex with respect to $\bm{P}$. Therefore, obtaining its optimal solution is challenging. In the subsequent section, we will present our developed time-efficient power allocation algorithm with property assurance.
@@ IV-A Algorithm Design
By following similar procedures, we can determine the required power for user $o_{M-2}$ to user $o_{1}$, which are served by waveguide $n$. Similarly, the transmit power for all other users can also be calculated. For each waveguide, based on all its served users’ transmit power, it is straightforward to obtain $P_{n}$ and $\bm{\alpha}_{n}$. However, this is not the final step. The update of the transmission power for other users with different waveguides will affect the decoding order of the NOMA users within this waveguide. This, in turn, impacts the power allocation among the NOMA users. Therefore, we need to repeat the previous steps iteratively until no user’s transmission power changes anymore. For brevity, we summarize the proposed power allocation method in Algorithm 1.
@@
8:  return  $\bm{\alpha}$ and $\bm{P}$
@@ IV-B Property Analysis
For brevity, we define $\bm{P}_{\bar{o}_{m}}$ as the transmit power of all users except user $o_{m}$ with waveguide $n$. It is evident that $p_{n,o_{m}}$ is a function of $\bm{P}_{\bar{o}_{m}}$, denoted as $p_{n,o_{m}}=f(\bm{P}_{\bar{o}_{m}})$, which we termed as the interference function of this user. Based on the discussion in Section IV-A, we know that for user $o_{m}$ with waveguide $n$, the required minimum transmit power is given by
$\begin{split}p_{n,o_{m}}&=f(\bm{P}_{\bar{o}_{m}})\\ &=(2^{\frac{\bar{R}_{n,o_{m}}}{W}}-1)(\sum_{j=m+1}^{M}p_{n,o_{j}}+\text{IIN}_{n,o_{m}}),\end{split}$ (11)
@@
According to (11), we can check that the interference function $f$ satisfies the following three properties: (1) $f(\bm{P}_{\bar{o}_{m}})\geq 0$; (2) if $\bm{P}_{\bar{o}_{m}}\geq\bm{P}^{\prime}_{\bar{o}_{m}}$, then $f(\bm{P}_{\bar{o}_{m}})\geq f(\bm{P}^{\prime}_{\bar{o}_{m}})$; and (3) for all $\beta>0$, $\beta f(\bm{P}_{\bar{o}_{m}})\geq f(\beta\bm{P}_{\bar{o}_{m}})$. If a function satisfies the three criteria, we say that it is standard. Based on [13], the iterative algorithm using a standard function converges to the unique fixed point. This completes the proof. ∎
@@ V Numerical Results
Fig. 2 demonstrates the convergence performance of the proposed power control algorithm for different configurations with the number of waveguides $N=2$ and $N=3$, as well as the transmission rates constraints $\bar{R}_{n,m}$ of $10$ Mbps, $15$ Mbps, and $20$ Mbps. Given the same initial conditions, the proposed algorithm successfully converges within merely ten iteration cycles, which demonstrates high convergence efficiency. Nevertheless, the convergence efficiency of our proposed algorithm is influenced by the transmission rates. Higher transmission rates necessitate higher power levels, leading to increased interference power during iteration, which consequently reduces the convergence efficiency. Furthermore, adding a waveguide to meet the transmission rates constraints of $10$ Mbps, $15$ Mbps, and $20$ Mbps requires additional transmit power level ratios of $54.7\%$, $74.0\%$, and $95.0\%$ respectively, which aligns with a $50\%$ increase in users and sum rate of the system. However, it is evident that the transmit power ratios increase with higher transmission rate constraints. As discussed previously, more stringent requirements of the transmission rate inevitably result in increased power interference among the users on different waveguides, which consequently compromises the system’s performance. Hence, the interval of the waveguides and pinches should be adequately maintained to achieve low power consumption under high transmission rate constraints.
@@
Fig. 3 compares the minimal transmit power between the proposed power control algorithm and the equal power allocation algorithm. Taking the case of $N=2$ and $R=10$ Mbps as an example, our proposed algorithm reduces the minimal transmit power by $65.8\%$ and $89.2\%$ for $M=2$ in Fig. 3a and $M=3$ in Fig. 3b, respectively, compared to the equal power allocation algorithm. The better performance of our proposed algorithm results from the fact that the upper bound of the transmission rate $\bar{R}_{n,o_{m}}\leq R_{n,o_{m}}\leq W\log_{2}(1+\frac{1}{M-o_{m}})$ is unavoidable in the equal power allocation algorithm, which can be proved through (7) and (5). Consequently, under high transmission rate constraints, the equal power strategy obtains the suboptimal solution that are either excessively large or computationally infeasible to obtain, which can be observed from the case of $\bar{R}_{n,m}=10$ Mbps and $\bar{R}_{n,m}=15$ Mbps in Fig. 3b. Furthermore, it is also illustrated in Fig. 3a that the minimal transmit power increases with the number of waveguides. Specifically, in systems with few waveguides, adding new waveguides requires significantly more power, whereas the pinching antennas system with more waveguides demands nearly equal power when adding new waveguides, since the slope remains almost identical over the interval $N\in[4,7]$ in Fig. 3a.
(a) $M=2$
(b) $M=3$
@@
Fig. 4 is displayed to reveal the relationship between the minimal transmit power and the interval of the waveguides and pinches. It is demonstrated that the minimal transmit power, as a function of the interval $D$, displays an approximate oscillatory decay with a negative exponential envelope. The observed functional behavior motivates us to approximate it with the underdamped sine function in relation to the interval $D$. More specifically, within the long ranges of $D\in[1,40]$, the function exhibits decay with a negative exponential envelope in Fig. 4a and Fig. 4b, while within the short ranges of $D\in[37,40]$, the function demonstrates periodicity in Fig. 4c and Fig. 4d. The fundamental reason lies in that the definition of channel gain $h_{m,n}$ in (2) can be represented by using the underdamped sine function. Therefore, the properties of the $\text{SINR}_{n,o_{m}}$ in (5), even and the minimal transmit power, share similarities with the underdamped sine function. To precisely reveal the insights in Fig. 4, we take the case of $R=10$ Mbps in Fig. 4b as an example. Specifically, each lower value of the minimal transmit power, i.e., $D=6$ meters, and $D=9$ meters, essentially indicates the better the channel condition, where the channel gain magnitude $|h_{n,m}|$ of the current waveguide significantly exceeds that of other waveguides. As the interval $D$ increases, the minimal transmit power approaches a constant. Since the increment $|\tilde{\psi}_{n,m}-\psi_{n,i}|=|m-i|D+d$ increases, thereby leading to a corresponding decrease in the channel gain magnitude $|h_{n,m}|$. Finally, more power must be allocated to meet the constraints of the transmission rates. Besides, given the large interval $D$, it follows that the increment $|\tilde{\psi}_{n,m}-\psi_{n,i}|=|m-i|D+d$ is the much larger coefficient for $\forall m\neq i$, compared to the increment $|\tilde{\psi}_{n,m}-\psi_{n,m}|=d$. Hence, the channel gain $h_{n,m}$ associated with user $m$ on waveguide $n$ is nearly independent of pinches far from the user. In other words, the $N$ correlated waveguides are decoupled into $N$ independent waveguides, and each channel gain magnitude $|h_{n,m}|$ is approximately expressed as $|h_{n,m}|\approx\sqrt{\eta}/d$. Evidently, the minimal transmit power of the NOMA-assisted pinching antenna systems tends to a constant as $D\to+\infty$. Besides, the decay rate and the degree of oscillation in Fig. 4 are influenced by the transmission rate constraints, which implies the tradeoff between transmission rate and NOMA-assisted pinching antenna system stability. As the transmission rate increases, the interference among users on other waveguides undoubtedly becomes more severe. Therefore, the design of the interval $D$ plays a crucial role in determining the minimal transmit power allocation of the system.

### v2

@@ II System Model
$\boldsymbol{x}_{n}=\sum_{m=1}^{M}\sqrt{P_{n}\alpha_{n,m}}\boldsymbol{x}_{n,m},$ (1)
where $\boldsymbol{x}_{n,m}$ and $\alpha_{n,m}$ denote the signal and the power allocation coefficient for user $m$ served by waveguide $n$, respectively.
@@
$h_{n,m}=\sum_{i=1}^{M}\frac{\sqrt{\eta}e^{-j\frac{2\pi}{\lambda}\lVert\tilde{\psi}_{n,m}-\psi_{n,i}\rVert}}{\lVert\tilde{\psi}_{n,m}-\psi_{n,i}\rVert}e^{-j\theta_{n,i}},$ (2)
[…]
$h_{n^{\prime},n,i,m}=\frac{\sqrt{\eta}e^{-j2\pi(\frac{1}{\lambda}\lVert\tilde{\psi}_{n,m}-\psi_{n^{\prime},i}\rVert+\frac{1}{\lambda_{g}}\lVert\psi_{n^{\prime},0}-\psi_{n^{\prime},i}\rVert)}}{\lVert\tilde{\psi}_{n,m}-\psi_{n^{\prime},i}\rVert}.$ (3)
@@
in which $\sigma^{2}$ represents the additive white Gaussian noise power. For the users served by the waveguide $n$, we rank their nSINRs in ascending order. Denote by $\boldsymbol{o}_{n}=(o_{1},o_{2},\ldots,o_{M})$, which satisfies $\text{nSIINR}_{n,o_{i}}<\text{nSIINR}_{n,o_{j}}$ if $i<j$. Following the downlink NOMA decoding principle, the interference plus noise ratio of user $o_{m}$ served by the waveguide $n$, denoted by $\text{SINR}_{n,o_{m}}$, can be expressed as follows:
@@ III Problem Formulation
In this paper, our objective is to minimize the total transmit power of all users while satisfying the data rate requirement of each user. For simplicity, we define $\boldsymbol{\alpha}_{n}=(\alpha_{n,1},\alpha_{n,2},\ldots,\alpha_{n,M})$ as the power allocation coefficients for users associated with the waveguide $n$. In addition, let $\boldsymbol{\alpha}=(\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2},\ldots,\boldsymbol{\alpha}_{N})$ be the power control coefficients for all users. We also denote $\boldsymbol{P}=(P_{1},P_{2},\ldots,P_{N})$ as the power allocation vector of all the waveguides. Furthermore, denote by $\bar{R}_{n,m}$ the minimum data rate requirement for user $m$ served by the waveguide $n$. Mathematically, we have
$\underset{\boldsymbol{\alpha},\boldsymbol{P}}{\text{min}}\sum_{n=1}^{N}P_{n}$ (8)
@@
where C1 indicates that, for each waveguide, the sum of the power coefficients of its served users should be 1. C2 ensures that the power allocation coefficient for each user lies between 0 and 1. C3 limits the maximum transmission power of waveguide $n$ to $\bar{P}_{n}$. Finally, C4 ensures that each user achieves its minimum data rate. The formulated optimization problem (8) is non-convex because $R_{n,m}$ is non-convex with respect to $\boldsymbol{P}$. Therefore, obtaining its optimal solution is challenging. In the subsequent section, we will present our developed time-efficient power allocation algorithm with property assurance.
@@ IV-A Algorithm Design
By following similar procedures, we can determine the required power for user $o_{M-2}$ to user $o_{1}$, which are served by waveguide $n$. Similarly, the transmit power for all other users can also be calculated. For each waveguide, based on all its served users’ transmit power, it is straightforward to obtain $P_{n}$ and $\boldsymbol{\alpha}_{n}$. However, this is not the final step. The update of the transmission power for other users with different waveguides will affect the decoding order of the NOMA users within this waveguide. This, in turn, impacts the power allocation among the NOMA users. Therefore, we need to repeat the previous steps iteratively until no user’s transmission power changes anymore. For brevity, we summarize the proposed power allocation method in Algorithm 1.
@@
8:  return  $\boldsymbol{\alpha}$ and $\boldsymbol{P}$
@@ IV-B Property Analysis
For brevity, we define $\boldsymbol{P}_{\bar{o}_{m}}$ as the transmit power of all users except user $o_{m}$ with waveguide $n$. It is evident that $p_{n,o_{m}}$ is a function of $\boldsymbol{P}_{\bar{o}_{m}}$, denoted as $p_{n,o_{m}}=f(\boldsymbol{P}_{\bar{o}_{m}})$, which we termed as the interference function of this user. Based on the discussion in Section IV-A, we know that for user $o_{m}$ with waveguide $n$, the required minimum transmit power is given by
$\begin{split}p_{n,o_{m}}&=f(\boldsymbol{P}_{\bar{o}_{m}})\\ &=(2^{\frac{\bar{R}_{n,o_{m}}}{W}}-1)(\sum_{j=m+1}^{M}p_{n,o_{j}}+\text{IIN}_{n,o_{m}}),\end{split}$ (11)
@@
According to (11), we can check that the interference function $f$ satisfies the following three properties: (1) $f(\boldsymbol{P}_{\bar{o}_{m}})\geq 0$; (2) if $\boldsymbol{P}_{\bar{o}_{m}}\geq\boldsymbol{P}^{\prime}_{\bar{o}_{m}}$, then $f(\boldsymbol{P}_{\bar{o}_{m}})\geq f(\boldsymbol{P}^{\prime}_{\bar{o}_{m}})$; and (3) for all $\beta>0$, $\beta f(\boldsymbol{P}_{\bar{o}_{m}})\geq f(\beta\boldsymbol{P}_{\bar{o}_{m}})$. If a function satisfies the three criteria, we say that it is standard. Based on [13], the iterative algorithm using a standard function converges to the unique fixed point. This completes the proof. ∎
@@ V Numerical Results
Fig. 2 demonstrates the convergence performance of the proposed power control algorithm for different configurations with the number of waveguides $N=2$ and $N=3$, as well as the transmission rates constraints $\bar{R}_{n,m}$ of $5$ Mbps, $10$ Mbps, and $15$ Mbps. Given the same initial conditions, the proposed algorithm successfully converges within merely ten iteration cycles, which demonstrates high convergence efficiency. Nevertheless, the convergence efficiency of our proposed algorithm is influenced by the transmission rates. Higher transmission rates necessitate higher power levels, leading to increased interference power during iteration, which consequently reduces the convergence efficiency. Furthermore, adding a waveguide to meet the transmission rates constraints of $5$ Mbps, $10$ Mbps, and $15$ Mbps requires additional transmit power level ratios of $51.5\%$, $57.3\%$, and $94.0\%$ respectively, which aligns with a $50\%$ increase in users and sum rate of the system. However, it is evident that the transmit power ratios increase with higher transmission rate constraints. As discussed previously, more stringent requirements of the transmission rate inevitably result in increased power interference among the users on different waveguides, which consequently compromises the system’s performance. Hence, the interval of the waveguides and pinches should be adequately maintained to achieve low power consumption under high transmission rate constraints.
@@
Fig. 3 compares the minimal transmit power between the proposed power control algorithm and the equal power allocation algorithm. Taking the case of $N=2$ and $R=10$ Mbps as an example, our proposed algorithm reduces the minimal transmit power by $76.0\%$ and $93.2\%$ for $M=2$ in Fig. 3(a) and $M=3$ in Fig. 3(b), respectively, compared to the equal power allocation algorithm. The better performance of our proposed algorithm results from the fact that the upper bound of the transmission rate $\bar{R}_{n,o_{m}}\leq R_{n,o_{m}}\leq W\log_{2}(1+\frac{1}{M-o_{m}})$ is unavoidable in the equal power allocation algorithm, which can be proved through (7) and (5). Consequently, under high transmission rate constraints, the equal power strategy obtains the suboptimal solution that are either excessively large or computationally infeasible to obtain, which can be observed from the case of $\bar{R}_{n,m}=15$ Mbps in Fig. 3(b). Furthermore, it is also illustrated in Fig. 3(a) that the minimal transmit power increases with the number of waveguides. Specifically, in systems with few waveguides, adding new waveguides requires significantly more power, whereas the pinching antennas system with more waveguides demands nearly equal power when adding new waveguides, since the slope remains almost identical over the interval $N\in[4,7]$ in Fig. 3(a).
@@
Fig. 4 is displayed to reveal the relationship between the minimal transmit power and the interval of the waveguides and pinches. It is demonstrated that the minimal transmit power, as a function of the interval $D$, displays an approximate oscillatory decay with a negative exponential envelope. The observed functional behavior motivates us to approximate it with the underdamped sine function in relation to the interval $D$. More specifically, within the long ranges of $D\in[1,40]$, the function exhibits decay with a negative exponential envelope in Fig. 4(a) and Fig. 4(b), while within the short ranges of $D\in[37.5,39]$, the function demonstrates approximate periodicity in Fig. 4(c) and Fig. 4(d). The fundamental reason lies in that the definition of channel gain $h_{m,n}$ in (2) can be represented by using the underdamped sine function. Therefore, the properties of the $\text{SINR}_{n,o_{m}}$ in (5), even and the minimal transmit power, share similarities with the underdamped sine function. To precisely reveal the insights in Fig. 4, we take the case of $R=10$ Mbps in Fig. 4(b) as an example. Specifically, each lower value of the minimal transmit power, i.e., $D=6$ meters, essentially indicates the better the channel condition, where the channel gain magnitude $|h_{n,m}|$ of the current waveguide significantly exceeds that of other waveguides. As the interval $D$ increases, the minimal transmit power approaches a constant. Since the increment $|\tilde{\psi}_{n,m}-\psi_{n,i}|=\sqrt{(m-i)^{2}D^{2}+d^{2}}$ increases, thereby leading to a corresponding decrease in the channel gain magnitude $|h_{n,m}|$. Finally, more power must be allocated to meet the constraints of the transmission rates. Besides, given the large interval $D$, it follows that the increment $|\tilde{\psi}_{n,m}-\psi_{n,i}|=\sqrt{(m-i)^{2}D^{2}+d^{2}}$ is the much larger coefficient for $\forall m\neq i$, compared to the increment $|\tilde{\psi}_{n,m}-\psi_{n,m}|=d$. Hence, the channel gain $h_{n,m}$ associated with user $m$ on waveguide $n$ is nearly independent of pinches far from the user. In other words, the $N$ correlated waveguides are decoupled into $N$ independent waveguides, and each channel gain magnitude $|h_{n,m}|$ is approximately expressed as $|h_{n,m}|\approx\sqrt{\eta}/d$. Evidently, the minimal transmit power of the NOMA-assisted pinching antenna systems tends to a constant as $D\to+\infty$. Besides, the decay rate and the degree of oscillation in Fig. 4 are influenced by the transmission rate constraints, which implies the tradeoff between transmission rate and NOMA-assisted pinching antenna system stability. As the transmission rate increases, the interference among users on other waveguides undoubtedly becomes more severe. Therefore, the design of the interval $D$ plays a crucial role in determining the minimal transmit power allocation of the system.